OpenStudy (anonymous):

Can someone please help me with the following question I have to submit before tomorrow. Suppose I is the 3x3 identity matrix and A and B are 3x3 matrices such that: E1E2A = B Where: E1is the elementary matrix obtained by performing -0.5 (Row 2) gives(=) Row 2 on I, and E2 is the elementary matrix obtained by performing Row 1 + 0.5 Row 2 gives(=) Row 1 on I. a) Write down E1 and E2 b) Write down E1 inverse and E2 inverse

OpenStudy (turingtest):

not sure what you mean by "gives (=)"

OpenStudy (unklerhaukus):

\[\rightarrow\]

OpenStudy (anonymous):

sorry ignore the word gives then... Just =. or you may even see it as a right arrow.

OpenStudy (turingtest):

ah so all you need to do is take the identity matrix and perform the given row operation to find En

OpenStudy (unklerhaukus):

\[\textbf I=\left[\begin{array}\ 1&0&0 \\0&1&0\\0&0&1\end{array}\right]\]

OpenStudy (anonymous):

I'm not completely catching on... I have I as above yes...

OpenStudy (anonymous):

does E1 then = \[\left[\begin{matrix}1 & 0 & 0 \\ 0 & 2 & 0 \\ 0 & 0 & 1\end{matrix}\right]\]

OpenStudy (unklerhaukus):

your missing the negative

OpenStudy (anonymous):

I see, Thus I must make the 2 a -2 in order for it to give poitive 1 when multiplied with -0.5?

OpenStudy (unklerhaukus):

\[\textbf I=\left[\begin{array}\ 1&0&0 \\0&1&0\\0&0&1\end{array}\right]\] \[\sim \left[\begin{array}\ 1&0&0 \\0&-\frac12&0\\0&0&1\end{array}\right] \qquad -\frac{R_2}2 \longrightarrow R_2\]

OpenStudy (anonymous):

And E2 = \[\left[\begin{matrix}1 & 0 & 0 \\ 0 & 0 & 0 \\ 0 & 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

Is that the correct final matrix of what E1 should be? Bit confused now.

OpenStudy (unklerhaukus):

\[\textbf I=\left[\begin{array}\ 1&0&0 \\0&1&0\\0&0&1\end{array}\right]\] \[\sim \left[\begin{array}\ 1&\frac 32&0 \\0&1&0\\0&0&1\end{array}\right] \qquad {R_1}+\frac {R_2}2 \longrightarrow R_1\]

OpenStudy (phi):

I assume that 3/2 in position 1,2 should be 1/2

OpenStudy (anonymous):

Yes because I find the problem there... Let me check

OpenStudy (anonymous):

Phi do you feel that this is the correct answer as to what matrix E2 should be?

OpenStudy (anonymous):

That row operations according to me doesn't give R1 on I

OpenStudy (phi):

the E2 matrix is I with the 2,1 position replaced with 1/2. If you multiply E2 times any 3x3 matrix it will add 1/2 row2 to row1 and put the answer in row 1 the other 2 rows of the matrix will be copied over. when they say E2 is the elementary matrix obtained by performing Row 1 + 0.5 Row 2 gives(=) Row 1 on I. they must mean when E2 is applied to matrix A ?

OpenStudy (unklerhaukus):

oh yeah 0 isn't 1 my mistake

OpenStudy (anonymous):

Thus E1 = \[\left[\begin{matrix}1 & 0 & 0 \\ 0 & -2 & 0 \\ 0 & 0& 1\end{matrix}\right]\]

OpenStudy (anonymous):

And E2=

OpenStudy (phi):

you mean -1/2 for E1 and -2 for the inverse of E1, right?

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 0.5 & 0 \\ 0 & 1 & 0\\ 0 & 0& 1\end{matrix}\right]\]

OpenStudy (anonymous):

Sorry looks like I still miss something apparently...:(

OpenStudy (phi):

that is E2. For the inverse of E2, negate the "key entry" (the 0.5)

OpenStudy (phi):

E1is the elementary matrix obtained by performing -0.5 (Row 2) that is -0.5 not -2

OpenStudy (anonymous):

-0.5 x (Row 2 of which matrix) to give what value?

OpenStudy (phi):

Here is an example of what is going on.

OpenStudy (phi):

Is there more to this question? Do they want you to find an A and B ?

OpenStudy (anonymous):

Yes : Write A as a product of E1 inverse, E2 inverse and B. Then determine A if B = \[\left[\begin{matrix}3 & 2 & 1 \\ -2 & 0 & 1 \\ 3 &1 & 1\end{matrix}\right]\]

OpenStudy (phi):

OK, everything makes sense except where they say "gives(=) Row 2 on I " the elementary matrix is defined by the row operations (not by the results) but you have E1 and E2 and can check that they do the specified row operation. with B= E1*E2*A then A= E2i*E1i*B (E1i is E1 inverse)

OpenStudy (anonymous):

When they ask me to Write A as a product of E1 inverse, E2 inverse and B. Must I multiply all 3 these matrices to obtain 1 matrix (the product)?

OpenStudy (phi):

I assume they want what I just posted, and then use that "formula" by plugging in the known B to find A

OpenStudy (anonymous):

Thank you so much, I'm completing the question now, if I wonder about anything still, I will ask. Thank you for all the help!

OpenStudy (phi):

yw

OpenStudy (anonymous):

phi, I get the matrix A as \[\left[\begin{matrix}4 & 2 & 0.5 \\ 4 & 0 & -2\\ 3 & 1 & 1\end{matrix}\right]\] is this correct?

OpenStudy (phi):

No, It looks like you applied the inv(E)'s in the wrong order. You were told \[ E_1E_2A = B \] left multiply both sides by inverse(E1): \[ E_1^{-1}E_1E_2A = E_1^{-1}B \] the left side simplifies to \[ E_2A = E_1^{-1}B \] now left multiply by the inverse of E2 \[E_2^{-1} E_2A = E_2^{-1} E_1^{-1}B \] or \[A = E_2^{-1} E_1^{-1}B \] as a check, you should get B back using \[ B=E_1E_2A \]